OpenStudy (anonymous):

The following Geometric Sequence is Given: 2, 2^(x+1), 2^(2x+1), 2^(3x+1),... i need to find 3 things. 1) the nth term An. 2)the fifth term A5. 3) the eight term A8. in order to find A5 and A8 i need to find the common ratio. The formula to find the common ratio is : r = a(k+1)/ak so far i have r = a2/a1 = 2^(x+10) / 2 my problem is that i don't know how to solve r = 2^(x+1) / 2

OpenStudy (akashdeepdeb):

Alright, so we have the sequence given as:\[2,~~2^{x+1},~~2^{2x+1},~~2^{3x+1}~~...\] Which can also be written as: \[2,~~2^x.2,~~2^{2x}.2,~~2^{3x}.2~~...\] Which again can be written as: \[2,~~~(2^x).2,~~~{(2^x)}^2.2,~~~{(2^x)}^3.2~~...\] NOTE: \(2^{ax} = {(2^x)}^a = {(2^a)}^x\) Now, this should give you a sort of intuitive feel about the common ratio. What is being multiplied to the previous term to get the next term? It is: \(2^x\) Which can also be calculated using the formula: \[r = \frac{a_{n}}{a_{n-1}} = \frac{a_{n+1}}{a_{n}}\] Now you have to calculate:\[a_n = a.r^{n-1}\] Can you do that now? :)

OpenStudy (anonymous):

How did you get \[2^x.2\] from \[2^{x+1}\]?

OpenStudy (akashdeepdeb):

That's what it means. Using the law of indices. Let's say we have: \[a^2.a^3\] This equals: \[a^{2+3} = a^5\]